HELIUM NANODROPLET ISOLATION SPECTROSCOPY USING AN EXTERNAL CAVITY QUANTUM CASCADE LASER

\underline{ALEX MORRISON} and GARY E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602.

A mode-hop free tunable external cavity quantum cascade laser (Daylight Solutions) is used to probe the vibrational dynamics of helium solvated H_{2}O and NO_{2} monomers and dimers. The demonstrated high output power and mode-hop free tuning of the laser in combination with a multipass cell and bolometric detection allow for high resolution spectroscopy studies of helium solvated molecules between 10 and 5 μm.